### v1

# Through the Lens of Split Vote: Exploring Disagreement, Difficulty and Calibration in Legal Case Outcome Classification

Abstract
In legal decisions, split votes (SV) occur when judges cannot reach a unanimous decision, posing a difficulty for lawyers who must navigate diverse legal arguments and opinions. In high-stakes domains, understanding the alignment of perceived difficulty between humans and AI systems is crucial to build trust. However, existing NLP calibration methods focus on a classifier’s awareness of predictive performance, measured against the human majority class, overlooking inherent human label variation (HLV). This paper explores split votes as naturally observable human disagreement and value pluralism. We collect judges’ vote distributions from the European Court of Human Rights (ECHR), and present SV-ECHR [^1] Our dataset and code will be available upon publication. a case outcome classification (COC) dataset with SV information. We build a taxonomy of disagreement with SV-specific subcategories. We further assess the alignment of perceived difficulty between models and humans, as well as confidence- and human-calibration of COC models. We observe limited alignment with the judge vote distribution. To our knowledge, this is the first systematic exploration of calibration to human judgements in legal NLP. Our study underscores the necessity for further research on measuring and enhancing model calibration considering HLV in legal decision tasks.
[…]
Through the Lens of Split Vote: Exploring Disagreement, Difficulty and Calibration in Legal Case Outcome Classification
[…]
1 Introduction
Figure 1: Split Votes in ECtHR Decisions
The task of Case Outcome Classification (COC) involves the classification of legal case outcomes based on textual descriptions of their facts. While achieving high performance is desirable, in high-stakes domains such as legal and medical decisions, the quantification of a model’s predictive confidence, or conversely, its uncertainty, is particularly valuable. The incorporation of uncertainty estimation allows experts to make more informed decisions, especially when the model may be uncertain or where the consequences of a misdiagnosis are significant. Evaluating whether models are aware of their limitations is known as calibration Guo et al. (2017); Desai and Durrett (2020). It assesses the extent to which predictive probabilities accurately reflect the likelihood of a prediction being correct. Models can opt to abstain when the uncertainty exceeds a predefined threshold — a method commonly referred to as selective classification El-Yaniv et al. (2010); Geifman and El-Yaniv (2017).
Current NLP research focuses on prediction confidence and calibration to assess a classifier’s awareness of its predictive performance only. This evaluation is commonly conducted against the human majority class. However, recent developments in NLP research have shed light on the prevalence of inherent human label variation (HLV) Plank (2022), observing disagreement [^2] The term HLV embraces disagreement and plausible variation. In this paper, we use these terms interchangeably. across various tasks Uma et al. (2021). Scholars in the field argue for the acknowledgment and acceptance of HLV, as it mirrors the diverse and pluralistic nature of human values Sorensen et al. (2024). Notably, Baan et al. 2022 has demonstrated that widely-used calibration metrics may not be suitable for datasets characterized by inherent human disagreement.
In light of these findings, we posit that models should not only be calibrated to recognize their own uncertainty (performance calibration) but also be equipped to discern instances where humans exhibit uncertainty (human calibration) Baan et al. (2022). This dual focus aims to foster trust among end-users and mitigates potential harm caused by models. Consequently, a critical aspect of this trust involves ensuring the alignment of perceived difficulty between human and models.
This motivates us to study split votes (SV) in court decisions (Fig 1). The judge vote ratio is a naturally occurring human disagreement at the European Court of Human Rights (ECtHR). We present SV-ECHR, a COC dataset with judge split vote information. We study the disagreement sources among judges from a legal, linguistic, and NLP perspective. We adapt and analyse the task-agnostic taxonomy components of Xu et al. 2023b and introduce SV-specific subcategories. We also quantitatively assess the effects of different subcategories on judges’ agreement using proxy variables. The results suggest that disagreements are mainly due to the social-political context of cases.
In addition, we assess the alignment of perceived difficulty between COC models and humans, as well as confidence- and human-calibration. While we see acceptable performance in perceived difficulty and confidence calibration, our analysis indicates suboptimal human calibration. This underscores the necessity for a more in-depth inquiry into methods to better align models’ calibration with human behaviour, highlighting opportunities for further research in this direction.
2 Related Work
COC has been referred to as Legal Judgment Prediction (LJP) in previous research Medvedeva et al. (2020). Under both terms, there exist numerous works involving decision corpora from the European Court of Human Rights (ECtHR) Aletras et al. (2016); Chalkidis et al. (2019); Tyss et al. (2023). All of these approaches mainly focus on the COC performance of models, which is commonly measured against the human majority class. Our SV-ECtHR dataset extends this work and contains the nuanced judges’ vote information. We proceed to use it to systematically investigate disagreement among judges and the alignment of perceived difficulty between models and judges.
Disagreement / Human Label Variation is receiving growing attention in mainstream NLP. Various works highlight the presence of HLV, emphasizing the abundance and plausibility of such human disagreements Pavlick and Kwiatkowski (2019); Plank (2022). Researchers advocate for embracing HLV for more trustworthy AI Talat et al. (2022); Casper et al. (2023). As real-world applications are used to assist diverse audiences, it becomes crucial to investigate and include pluralistic human values in NLP systems Sorensen et al. (2023). This motivates our study of disagreements among judges’ split-votes in the legal decision process. Recently, several studies have proposed task-specific taxonomies to identify potential sources of disagreement in various NLP tasks Uma et al. (2021); Sandri et al. (2023); Jiang and Marneffe (2022). Building on the meta-analysis of these existing taxonomies, Xu et al. 2023b generalize them to two layers of task-agnostic categories and introduce task-specific categories for COC rationale annotation. In this work, we study judges’ disagreements in case decision votes, presenting a taxonomy of disagreement with SV-specific subcategories.
Perceived Difficulty Regarding model difficulty, there is increasing interest in identifying difficult data instances. Various techniques, such as Influence Functions Koh and Liang (2017) and training loss Han et al. (2018) have been proposed to identify the difficulty of data instances to a certain model. Pointwise V-usable information (PVI) is a recently introduced difficulty metric Ethayarajh et al. (2022), which incorporates mutual information and other types of informativenessXu et al. (2020). Despite its recent introduction, PVI has gained significant attention and proven effective in various tasks, including rationale evaluation Prasad et al. (2023) and data selection for augmentation Lin et al. (2023). In this work, we leverage PVI to evaluate the alignment of perceived difficulty between human and COC models.
Calibration Existing calibration studies primarily focus on a classifier’s awareness of its predictive performance, commonly evaluated against the human majority class Guo et al. (2017); Desai and Durrett (2020). Baan et al. 2022 argue that calibrating against the human majority class is not meaningful in settings with inherent human disagreement. In our work, we assess the performance and human calibration of COC models in the context of split-vote, a naturally occurring human disagreement. To the best of our knowledge, this is the first systematic exploration of human calibration in the legal domain.
3 Legal and Linguistic Backgrounds
The Legal Decision Making Process at the ECtHR begins with the applicants lodging their accusation, alleging one or more violations of articles of the European Convention on Human Rights (ECHR). This allegation process corresponds to Task B (allegation prediction) in the widely used LexGLUE benchmark Chalkidis et al. (2022a). After receiving the complaint, the court undertakes a review of the case, aiming to determine whether a violation has indeed occurred, corresponding to Task A (violation prediction) in LexGLUE. Cases falling under the purview of well-established ECtHR case-law are directed to a three-judge Committee, while others may find themselves before a seven-judge Chamber, where decisions are reached through a majority vote. In some circumstances, the Grand Chamber, comprising 17 judges, adjudicates cases referred to it by request. In the process of rendering a judgment, the Court typically examines only the specific articles alleged by the applicant. This procedure aligns with Task A|B (violation prediction given allegation), as introduced by Santosh et al. 2022. Consequently, our study evaluates the uncertainty of COC models for Task A|B, mirroring the real legal process. Sources of Disagreement in Legal Decisions Recent research has delved into sources of HLV and disagreement across various NLP tasks from a linguistic perspective. The classic framework “Triangle of Reference" Ogden and Richards (1923) has been widely adopted to categorize disagreement sources in classification tasks Aroyo and Welty (2015); Jiang and Marneffe (2022). This concept primarily addresses the relationship between linguistic symbols and the corresponding objects they represent. To study uncertainty sources in NLG tasks, Baan et al. 2023 extend it to the “Double Triangle of Language Production” catering to the complexities of language generation. However, the legal decision-making process introduces an additional layer of complexity. Legal scholars conceptualize the decision-making process in case-law as circular Ichim (2019). To better capture this nature, we propose the adoption of the "Direction of Fit" (DoF) framework from speech act theory Searle and Vanderveken (1985). In our proposed DoF framework for the legal decision-making process (Fig 4 in App A), judges render decisions based on the interpretation of the existing normative structure of case-law (Law-to-Case). Simultaneously, these case decisions serve as a foundation for future litigation (Case-to-Law). We extend this model with a temporal axis to capture the continuous reproduction of the main normative model over time.
Figure 2: Taxonomy of disagreement sources among judges in legal decisions and proxy labels.
4 The SV-ECHR Dataset
Dataset Collection We extract the judge votes distribution for each alleged article [^3] Following Chalkidis et al. 2022a; Santosh et al. 2022 We use only the 10 most prominent ECHR articles. from the public database HUDOC [^4] https://hudoc.echr.coe.int using regular expressions. The information about judge votes is always present in the conclusion section and generally follows certain patterns (See more details in App B). We did a two-round quality assessment. [^5] App C offers details on the quality assessment process. The F1 score of our regular expression rules increased from 0.81 in the first round to 0.98 in the second round.
Dataset Analysis: We then merge the vote distribution to the ECHR A|B dataset Santosh et al. (2022) according to the document ID. We name the merged dataset SV-ECHR. It consists of 11k case fact descriptions along with target label information about which convention articles have been alleged to be violated (task B), and which the court has eventually found to be violated (task A), and the judges vote distribution of each alleged article. The dataset is chronologically split into training (2001–2016), validation (2016–2017) and test set (2017-2019) with 9k, 1k, 1k cases respectively. The label set includes 10 prominent ECHR convention articles. On average, each case has around 1.6 alleged articles. Among all 17,604 alleged case-alleged article pairs (hereafter pair), only round 7% are split-voted. These numbers underscore the significant label imbalance within the dataset. Additional statistics can be found in Tab 5 in App D. To investigate the extent of judges’ disagreement in SV-ECHR, we assess the entropies of vote distributions for each pair. For the detailed calculation and histogram of the entropy distribution, see App F. We see a large share ($\sim$ 60% of chamber votes) of single dissenting votes, where only one single judge voted differently than the six judge majority. A similar pattern is observed for Grand Chamber cases involving 17 judges. This pattern aligns with other observations made in empirical legal scholarship (Fobbe 2022, see App G).
Correction of Inconsistent Metadata Xu et al. 2023b already pointed out inconsistent allegation information in HUDOC. During our quality control process, we found further inconsistent violation information in approximately 2% of the training set. We updated SV-ECHR with the correct metadata. [^6] See App E for more details of our metadata correction Our finding of such inconsistencies calls for mindful data curation when developing COC datasets.
5 Disagreement among Judges’ Votes
5.1 Disagreement Taxonomy
We utilize the disagreement taxonomy from Xu et al. 2023b to analyze the reasons behind judges’ split votes. Fig 2 displays the taxonomy, with two adapted task-agnostic levels and our expanded split-vote specific subcategories and proxy variables. [^7] For a comprehensive understanding of each taxonomy category, we direct the reader to Xu et al. 2023b In the following sections, we provide detailed explanations for each SV-specific category.
5.1.1 The Data
Disagreements among judges can stem from various case aspects, including Genuine Ambiguity within the Normative Structure, Narrative Complexity of the facts, and/or Specific Legal Context.
Genuine Ambiguity is attributed to Normative Uncertainty of case law in the context of legal NLP (Xu et al. 2023b), which emerges when the court is presented with the possibility of justifying an outcome through multiple legal source interpretations and argumentation. Its occurrence is not uncommon in ECtHR judgments due to the deliberate drafting of the convention in a ‘flexible’ manner to allow tailoring the interpretation to domestic specificities based on the subsidiarity and margin of appreciation principles. It should be noted that there exists and analog factual ambiguity where the facts of the case are unclear based on limited (or contradicting) evidence provided. As the ECtHR does not engage in evidentiary reasoning, this aspect is out of scope for this work.
Text Instantiation covers inconsistency, incompleteness, or biases during text production (i.e. judgment document drafting). When reviewing case files, judges analyze materials that encapsulate the factual background, legal arguments, and evidence presented by both parties. However, these documents are usually produced by the Court Registry, priming the language accordingly. The likelihood of encountering textual unclarity, such as inconsistent framing of facts, increases with the length and complexity of the docket. We refer to this challenge as Narrative Complexity.
Context Sensitivity refers to the Specific Legal Context in the legal domain, characterized by undefined and controversial social-political factors for which judges may not find a precisely tailored legal explanation. We use the term ‘Context Specificity’ because pinpointing the exact arrangement of factors (i.e., attitudinal, normative, and strategic) influencing context in a specific decision is challenging Shapiro (2002). Certain social and political factors may only become contentious in a particular context and arrangement of factors.
5.1.2 The Annotator/Judge
Disagreements can also arise due to variation in human behavior, which we systematize as Noise, Subjectivity and Longitudinal behavioural change. Noise covers errors due to annotator’s Sloppy Annotation or Interface Limitation Uma et al. (2021); Sandri et al. (2023). We expect negligence-related noise and interface limitations to be negligible in naturally occurring ECtHR judge votes. [^8] Empirical legal scholarship has been investigating the relationship between judge productivity and caseload, including its ramifications on decision quality (e.g., Engel and Weinshall 2020). To the best of our knowledge, no such evidence has been documented in the ECtHR context.
Subjectivity The ECtHR is composed of judges from the 46 member states with diverse legal traditions and cultural backgrounds. Previous work in general NLP demonstrated how annotators’ demographic identities can influence how they label toxicity in text Sap et al. (2022). The judges’ Cultural Background, stemming from their legal training background and political consideration from their native countries may lead to variations in their rulings compared to those of their colleagues Voeten (2008). [^9] See discussion about National Judge in App I Divergent opinions may arise when, for example, two judges from the same country prioritizing different values. Pluralistic Human Values have recently been made a primary object of mainstream NLP research Sorensen et al. (2024). Political science research has explored this in the context of quantitative justice ideal point estimation on the US Supreme Court (SCOTUS, typically a scalar dimension between liberal and conservative, see Segal and Cover 1989, Martin and Quinn 2002). More similar to our focus on agreement, Ruger et al. 2004 trained an ensemble of court-level and judge-specific decision trees, and observed different performance on conservative and liberal judges.
A comparable liberal vs conservative investigation in the ECtHR context would not be useful. It is composed of 46 Member States, each with its own spectrum of political ideologies, more pluralist and complicated than the liberal-conservative divide. We account for some such disagreement based on the democratic score of the State they are representing (the HighRep variable, see below). Moreover, there is no temporal stability in the judges’ behavior because they have a limited mandate and act in different bench formations (chamber, grand chamber, etc.). Focusing on individual ideologies to the exclusion of variables accounting for the wider court dynamics could not support accurate analysis. We have therefore chosen to focus instead on normative uncertainty and the specific legal context.
The Longitudinal Dimension in Xu et al. 2023b considers two aspects: Individual Behavioral Change over time and Case Law Change. The latter involves shifts in the collective societal attitude towards specific phenomena over time, including changes in laws, policies, and cultural norms. In the legal domain, legal professionals must continually adapt their knowledge, strategies and reasoning to align with the ongoing evolvement of jurisprudence and applicable statutory law. Due to the limited scope of this work, we defer a comprehensive longitudinal study to future research.
5.2 Conflation of Categories
It is important to acknowledge that our taxonomy and its categories are not always sharply delineated. Conflation can occur and the boundary between categories may be blurry. Some sources can be interpreted as either Law-to-Case or Case-to-Law. For instance, cases involving vulnerable applicants often involve complex social-political issues lacking clear legal precedent, causing disagreement on the bench. Simultaneously, the judges’ subjectivity may lead to varying opinions on whether an applicant should be deemed vulnerable, i.e. judge subjectivity and legal context interact.
5.3 Proxies & Disagreement Correlation
To quantitatively explore the influence of different taxonomic categories on judges’ votes, we work with a legal expert to identify proxy variables. We hypothesize that these proxies correspond to higher disagreement among judges and evaluate the statistical association between them and the entropy of the vote distribution. If not otherwise mentioned, we retrieve relevant information from the HUDOC dataset and label a case as 1 if it is listed with the proxy feature and 0 otherwise.
For Normative Uncertainty, we choose KeyCase as a proxy following Xu et al. 2023b. The court annually chooses a set of “key cases”, which often deal with complex and novel legal issues. Given the absence of established legal standards for them, they often generate controversy and disagreement among experts. We hypothesize that judge tend to disagree more in “key cases”.
Specific Legal Context via two proxy variables:
HighRepCountry (High Reputation Country): Previous legal and political science work indicates ECtHR Judges tend to have more split votes in cases where the defending country is regarded as High Reputation Country Dothan (2014) . [^10] Countries with higher democracy index. See App K.
VulnApplicant (Vulnerability of Applicant): the ECtHR court adapts convention standards to meet vulnerable individuals’ needs despite vulnerability not being defined in the Convention. It remains underdefined within the context of ECtHR because the judges do not explain the process though which they identify an individual as vulnerable Heri (2021), which can lead to divergent legal opinions. We retrieve relevant information from the VECHR dataset Xu et al. (2023a) and categorize a case as 1 if its applicant is regarded as vulnerable. [^11] Only annotation of cases under Article 3 are available in VECHR before 2016. Hence our t-test of VulnApplicant is only on cases of Article 3. App L offers more explanation on Vulnerability in ECtHR jurisprudence.
Subjectivity refers to the disagreement due to personal opinions and values. We hypothesize that the GrandChamber, which consists of 17 judges, tends to more disagreement than the 7-judge Chamber.
5.3.1 Do proxy measures correlate with judges’ votes?
To measure the influence of each selected taxonomy category, for each binary proxy variable, we compute the entropy of the vote distribution among all cases exhibiting that variable (‘present’, value 1) and those that do not (‘absent’, value 0). We perform an independent t-test to compare the mean entropies between the two groups.
mean entropy t-value p-value
Proxy (absent/present): 0 1
GrandChamber 0.50 0.49 1.61 0.11
HighRepCountry 0.48 0.51 -4.28 2e-5*
VulnApplicant 0.50 0.62 -1.8871 0.006*
KeyCase 0.50 0.48 2.46 0.002*
Table 1: Associations of proxy variables and entropy of the votes distribution ($\ast$: p < 0.05)
mic-F1 mac-F1 hm-F1
BERT 72.31 ± 4.09 65.53 ± 6.22 64.22 ± 3.46
LegalBERT 73.55 ± 0.62 68.75 ± 1.13 68.03 ± 1.29
Table 2: COC performance on test set
Tab 1 gives the results. It shows significantly lower agreement among judges’ votes for HighRepCountry and VulnApplicant. In other words, specific contexts of case facts correlate more with disagreement than Normative Uncertainty. Judges express disagreement related to how facts fit the norm because of circumstances related to the social standing of the applicant and the political situation of the defending State. This is an inherent characteristic of adjudication where judges qualify the context so as to fit the norm. The inclusion of a category related to the legal norm’s scope could have provided further insights.
Our hypothesis regarding GrandChamber does not hold precisely because it does not matter which bench formation deals with the case - Chamber or Grand Chamber. What matters is how judges decide to fit the facts to the norm.
Interestingly, we found less disagreement among judges on KeyCase. Experts suggest this may stem from the Court’s assigning the importance score ex post, after interpretation and agreement are reached. The Court is more likely to designate a consensus case as KeyCase to maintain coherent jurisprudence, rather than a controversial case that could invite future applicants to base cases on disputed facts and dissenting opinions. Our DoF framework (App A) suggests that Case to Law uncertainty inherently perpetuates Law to Case uncertainty.
6 Judge-Model Misalignment
hm-F1 $\uparrow$ ECE $\downarrow$ DistCE $\downarrow$ count
/ TS soft / TS soft / TS soft
[…]
Table 3: COC performance (hm-F1), confidence-calibration (ECE), and human-calibration (DistCE) performance with std ($\pm$) on test set. "/": COC finetuned models in § 6.1; "soft": models fine-tuned with soft-loss; "TS": models after Temperature Scaling. Results shown over 3 random seeds. See Tab 8 in App R for performance on dev set.
hm-F1 mean PVI t-value p-value
U. SV U. SV
B 66.19 ± 3.55 43.76 ± 4.23 0.94 0.06 3.12 0.002
L 69.3 ± 1.88 53.67 ± 4.59 0.77 -0.12 2.66 0.008
Table 4: COC prediction scores (hm-F1) and difficulty scores (mean PVI) for cases within the unanimous (U.) and SV categories. B: BERT; L: LegalBERT
6.1 Experimental Setup of COC
We assess the alignment of a model’s prediction difficulty [^12] Another approach would be assessing a model’s prediction uncertainty. Refer to App N for an explanation of our decision to abstain from using Uncertainty evaluation methods with judge disagreement, specifically COC models trained on ECtHR Task A|B. To address input length constraints, we employ a BERT variant of the hierarchical attention model [^13] See App O for model structure and implementation. Yang et al. (2016), as adapted from Santosh et al. 2022.
Task Benchmark task ECtHR A|B (Violation Identification given Allegation) identifies the set of violated convention article(s) from the textual case facts and list of allegedly violated articles.
Models Santosh et al. 2022 encodes allegation information as a multi-hot vector concatenated with the text representation. By contrast, in this study, we textify allegation information as “Alleged Article: X”, where is X is a comma-separated list of articles and prepend to the input text. To explore the impact of domain-specific pre-training on uncertainty representations, we employ two BERT variants as the backbone in hierarchical models: (i) BERT Devlin et al. (2019) bert-base-uncased, trained on general English; (ii) LegalBERT Chalkidis et al. (2020) nlpaueb/legal-bert-base-uncased pretrained on English legal text, including ECtHR.
Metrics Our evaluation of COC performance employs the micro-F1 (mic-F1), macro-F1 (mac-F1), and hard-macro-F1 (hm-F1; Santosh et al. 2022). The hm-F1 is calculated as the mean F1-score for each article, where cases with that article having been violated are considered as positive instances, and cases with that article being alleged but not found to be violated as negative instances, resulting in a smaller pool of more difficult negatives.
COC Performance: Tab 2 reports the results of classification performance. Notably, the model with legal-specific pre-training (LegalBERT) outperforms the one with general pre-training (BERT). Although the mic-F1 and mac-F1 performances closely align with those reported in Santosh et al. (2022); Xu et al. (2023b), differences of $\sim$ 2%, the hm-F1 performance demonstrate a noteworthy improvement of $\sim$10%. [^14] This can be partly attributed to our correction of labels (see § 4). Some variation may also be due to randomness in weight initialization. See our discussion in App M Importantly, hm-F1 exhibits particular sensitivity to allegation information. For the rest of this paper, we adopt hm-F1 as our primary metric for COC performance.
6.2 Difficulty of SV Cases in COC
Pointwise V-usable information (PVI; Ethayarajh et al., 2022) measures the difficulty of an instance within a dataset for a given model as: $PVI(x\rightarrow y)=-\log_{2}g[x](y)+\log_{2}g^{\prime}[\emptyset](y)$. Computing PVI involves fine-tuning a model $g$ on two datasets. The first dataset $D$ comprises input-target pairs $\{(x_{i},y_{i})|(x_{i},y_{i})\in D\}$, while the second $D^{\prime}$ is used to fine-tune model $g^{\prime}$ on null-target pairs, $\{(\emptyset,y_{i})|(x_{i},y_{i})\in D\}$ ($\emptyset$ represents an empty string intended to fit the label distribution). PVI serves as the measure of information gain resulting from the provision of an input during fine-tuning. Higher PVI suggests a better representation of the input in the model, and thus an easier instance.
Experiment Settings: We use the COC fine-tuned models from Sec 6.1 as the input-target model $g$. We fine-tune an architecturally identical model $g^{\prime}$ with the input replaced by $\emptyset$. Subsequently, we partition the test set into two distinct partitions of SV and unanimous cases, respectively. Our expectation is that the SV subset will exhibit lower average PVI compared to the unanimous subset. We use entropy of the judges’ vote distribution as an estimator of case difficulty: the more the judges disagree with each other, the harder a case is.
Results: We present the average COC performance (hm-F1) and perceived difficulty (mean PVI scores) of the models in Tab 4. Both models exhibit lower hm-F1 and PVI scores for SV cases compared to unanimous cases. An independent t-test on the average PVI scores between the two groups reveals a significant difference ($p<0.05$). This indicates that SV cases are more challenging for the models than unanimous cases, aligning with human perceptions of difficulty. We also calculate the Pearson correlation coefficients between PVI scores and the entropy of judges’ vote distribution of SV cases. For BERT, the correlation coefficient (r-value) is -0.089, with a p-value of 0.38. LegalBERT has a r-value of -0.068 and a p-value of 0.48. The results reveal a negative correlation between PVI scores and the degree of disagreement among judges, consistent with our expectations. However, the observed correlation is very weak ($|r|<0.1$) and statistically insignificant ($p>0.05$). This suggests that, while models can capture differences in difficulty between unanimous and SV cases, they struggle to accurately represent the nuanced degree of judges’ disagreement. Therefore, we next evaluate the models’ confidence and soft-label training.
6.3 Calibration of SV Cases
6.3.1 Evaluation Metrics
Figure 3: The distribution over instance-based absolute errors between probabilities for class 1 of the model vs human vote distributions. See Fig 8 in App R for more figures comparing human uncertainty to model uncertainty
Most current NLP research focuses on Confidence Calibration Jiang et al. (2021); Desai and Durrett (2020): The model should be unsure when it does not know the answer. A model is considered well confidence-calibrated if its prediction confidence aligns with its predictive accuracy, commonly evaluated against the human majority class. Yet there is a growing interest in accounting for human label variants and/or pluralistic values Plank (2022); Sorensen et al. (2023). Therefore we extend our evaluation to include Human Calibration Baan et al. (2022): The model should be unsure when humans are unsure about the answer. We consider a model well human-calibrated if the categorical distribution of predicted class probabilities align well with the actual human vote distribution.
6.3.2 Calibration Methods
For confidence calibration, we employ Temperature Scaling (TS; Guo et al. 2017). This simple yet widely used post-hoc method uses a single temperature parameter $t$ to scale the output logits of a classifier. We choose the temperature $t$ by searching a range of possible values for $t$ on the dev set. For human calibration, we adopt the approach from Peterson et al. 2019; Uma et al. 2020 by finetuning with a Soft Loss Function. [^15] See App P for details of the Soft Loss Function. During training, models are exposed to “soft labels” derived from the judges’ vote distributions, serving as target distributions in a cross-entropy loss function.
Expected Calibration Error (ECE) is the most often used metric for confidence calibration Naeini et al. (2015); Guo et al. (2017). A lower ECE indicates better calibration, suggesting that the model’s predicted probabilities are more accurate reflections of the true probabilities. Refer to the App Q for further details on ECE.
DistCE by Baan et al. 2022 measures human calibration, which can be calculated as $\mathrm{DistCE}(x)=1/2\|\mathbf{q}-\mathbf{p}\|$, where $q$ is the vote distribution and $p$ is the model’s predictive distribution.
6.3.3 Results & Discussion
Tab 3 presents the experiment results. We make the following observations:
COC Performance COC Performance: Tab 3 indicates no improvement in COC performance (hm-F1) for SV cases in the test set, although improvement is observed for the development set (Tab 8 in App R). The scarcity of SV instances aside, temporal shifts in the dataset, as discussed in the longitudinal category (§ 5.1.2), may contribute to this discrepancy. SV cases could stabilize after a key legal decision. In other words, something that is legally uncertain in the training set time period may be settled law in the test dataset partitions. This suggests potential overfitting of our models to the training set timeframe. We encourage future models to incorporate the temporal dimension during training to mitigate such issues.
Confidence Calibration The implementation of TS greatly reduces ECE, suggesting enhanced confidence calibration for both models. However, fine-tuning with soft-loss to human labels yields minimal ECE changes with a discrepancy: BERT shows a slight decrease, while LegalBERT displays a minor increase. This mirrors the issue identified by Baan et al. 2022, highlighting the challenge of using ECE to measure calibration on disagreement data. They argue that even a classifier perfectly modeling human judgment distribution would still be severely miscalibrated when measured by ECE.
Human Calibration Soft-loss tuning slightly improves the alignment between the model’s prediction confidence and the human vote distribution (lower DistCE). Soft LegalBERT exhibits noticeably fewer instances of over-confident predictions when judges do not unanimously agree, as depicted on near-0/1 probability portions of Fig 3b, in contrast to the original model shown in Fig 3a. TS models exhibit substantially lower DistCE scores on split-vote cases than soft models. This may be attributed to an overly aggressive temperature, resulting in a more uniform output distribution (as shown in Fig 3c, with a temperature of 5.5). Moreover, judge votes often exhibit a quasi-bimodal distribution, with many split votes caused by a single judge dissenting pro/con the finding of a violation (refer to our discussion on single dissenting votes in § 4). Therefore, while TS models may display lower DistCE scores than soft models, they are not a good fit. It is crucial to analyze the distributions before drawing conclusion.
7 Conclusion
We present SV-ECHR, a COC dataset enriched with split-vote information. Our analysis of judges’ vote distributions leads to a taxonomy of disagreement sources. Our experiments reveal critical issues with current calibration measures, such as using confidence calibration metric ECE on data with inherent human disagreement. Additionally, caution is necessary when interpreting ECE and DistCE scores for human calibration. We call on the community to explore methods for measuring and improving the alignment of model calibration with human behavior; as well as more research into incorporating human disagreement and variability in NLP.
Limitations
Our study is constrained by the datasets, models, and selective prediction techniques under consideration, primarily relying on the ECtHR dataset. Expanding the investigation to encompass diverse datasets and legal jurisdictions would enhance our understanding of disagreement in judge decision votes and the alignment of perceived difficulty between judges and models.
Additionally, due to computational limitations, we are constrained from pre-training language models from scratch or fine-tuning large language models (LLMs). Our study relies on existing pre-trained BERT-based models, focusing solely on fine-tuning. We refrain from exploring LLM models, as no widely agreed-upon method for measuring calibration for LLMs has emerged at the time of submission. Furthermore, with respect to variations introduced by prompts and data contamination during pretraining, exploring the use of LLMs for difficulty perception and calibration on a small-scale, specialized legal dataset is a distinct research question deserving a separate paper.
Ethics Statement
In this study, our retrieved judges’ vote information from the publicly available HUDOC dataset, with the overarching goal of improving the alignment of perceived difficulty and calibration between models and judges. While these decisions include real names and are not anonymized, we do not anticipate any harm beyond the availability of this information resulting from our experiments.
The task of case outcome classification raises significant ethical and legal concerns, both generally and specifically concerning the European Court of Human Rights Medvedeva et al. (2020). It is important to clarify that we do not advocate for the practical implementation of COC within courts. Previous work Santosh et al. (2022) has demonstrated that these systems heavily rely on superficial and statistically predictive signals lacking legal relevance. This underscores the potential risks associated with employing predictive systems in critical domains like law and highlights the importance of trustworthy and explainable legal NLP.
In this work, we investigate the sources of disagreements among judges in their decisions on case facts. While technically situated outcome classification models, we intend our analysis of different types of disagreement to promote the acceptance of human label variation and pluralistic human values within the legal NLP community. By acknowledging and understanding various perspectives, interpretations, and biases of judges, we contribute to a more comprehensive and inclusive discourse within the field. A continuation of this work can unfold long term practical implications: By identifying patterns which create uncertainty, applicants could potentially ‘exploit’ the distinctive circumstances that change the normative assessment for the purpose of their own case, or otherwise inform litigation strategy. By the same token, however, it would offer the court the possibility to investigate and ‘check’ whether it applies legal norms coherently, in line with the demands of consistency, foreseeability and certainty. The potential effects of data-driven decision making in the legal domain cut both ways, and must be reconciled mindfully.
Appendix A DoF of Sources of Disagreement in Legal Decision Process
To better capture the circular nature of legal decision process, we propose the adoption of the "Direction of Fit" (DoF) framework from speech act theory Searle and Vanderveken (1985). In our proposed DoF framework for the legal decision-making process (Fig 4), judges render decisions based on the interpretation of the existing normative structure of case-law (Law-to-Case). Simultaneously, these case decisions serve as a foundation for future litigation (Case-to-Law). We extend this model with a temporal axis to capture the recurrent nature of this process and the continuous reproduction of the main normative model over time.
Figure 4: Schematic Diagram of Sources of Disagreement in Legal Decision Process
Appendix B Information about Judge’ Votes in HUDOC
We observe that the information about judge’ votes is always present in conclusion section of the decision towards the end and generally following certain patterns, such as: "Holds, by X votes to Y, that there has been a/no violation of Article Z of the Convention"; or "Holds by X votes to Y that Article Z has (not) been violated" as in Fig 6.
Appendix C Quality Validation of SV-ECHR
To ensure the quality of the dataset, we did a two-round quality assessment. In the first round, we manually retrieved the split-vote information of 20 cases as gold labels. We then assess the votes information exacted by our regex rule with the gold labels. We analysis the error and improve our regex rule. In the second round, we repeat this process with another 20 cases. The F1 score increased from 0.81 in the first round to 0.98 in the second round.
Appendix D Dataset Statistics
Tab 5 offers additional statistics on the SV-ECHR dataset.
# Cases
# Case
-Article Pairs
# alleged Case-
Article Pairs
# SV Case-
Article Pairs
Train 9000 90000 14513 960
Dev 1000 10000 1516 135
Test 1000 10000 1575 112
Table 5: More statistics of the SV-ECHR dataset.
Appendix E Correction of Inconsistent Metadata
# Case-Article Pairs
# alleged
pairs
# pair w/.
wrong alleg.
# pair w/.
wrong vio.
Train 90000 14513 2719 292
Dev 10000 1516 247 0
Test 10000 1575 278 0
Table 6: Statistics about the corrected meta information.
Tab 6 shows the statistics about our correction of meta information. Notably, apart from the inconsistent allegation information pointed out by Xu et al., 2023b. We further found approximately 2% inconsistent violation information in the train set during our quality control process. Most of the cases with so-called striked-out cases which are difficult to parse. We also find finetuning on the corrected train set improves model’s COC performance (§ 6.1) as discussed in App M.
Appendix F Entropy of Judges’ Vote Distribution
Fig 5 display the histogram of the entropy of each pair’s vote distribution. The entropy is calculated with $\mathbf{H}(\mathbf{p})=-\sum_{i\in\mathcal{C}}p_{i}\log\left(p_{i}\right)$ and $p_{i}=\frac{n_{i}}{\sum_{j\in\mathcal{C}}n_{j}}$ , where C is the label category set [violation, non-violation] and $n_{i}$ is the number of judge voting for category i.
Appendix G Debates in Legal Scholarship about Single Judge Dissenting Vote
Tab 5 reveals that $\sim$ 60% of chamber votes (involving 7 judges) exhibit entropy of around 0.4, indicating that one judge vote differently from the remaining 6 judges (entropy([1,6]) $\approx$ 0.4). A similar pattern is observed for Grand Chamber cases involving 17 judges. This Single Judge pattern aligns with ongoing debates in legal scholarship. For instance, Fobbe 2022 examined the number of dissenting opinions in decisions from the International Court of Justice (ICJ). Their results indicate a significant proportion of unanimous decisions, followed by a monotonously decreasing number of dissents. Some legal scholars support the view that diverging opinions do not signal a division of the bench inside the Court over the scope of protection of rights, even less a lowering of standards of protection. While other scholars, support the contrary view according to which dissenting opinions send a signal of walking-back in terms of effective protection. Tab 5 and Fobbe 2022’s finding actually shows that the case law does not consistently but rather seldomly give rise to dissenting opinions. Moreover, one can not ignore the National Judge’s dissenting vote (App I). It is difficult to assess what is the weight of a single judge dissenting on how the court will decide cases. Often times, dissenting opinions are not followed upon. What is of relevance is to determine a yardstick against which to measure disagreement in human decisions and this is what we are doing.
Figure 5: Histogram of entropy of judges’ votes distribution over train/dev/test split
Figure 6: Screenshot of Judges’ vote information in HUDOC
Appendix H Judge’s Negligence Errors
Empirical legal scholarship identified time constraints leading to judge negligence errors in certain jurisdictions like the US Board of Veterans’ Appeals Ho et al. (2019) and Israeli Magistrate Courts Engel and Weinshall (2020). However, to the best our knowledge, there is no evidence has been documented in the ECtHR context to support judges’ careless behavior. Further, the Hungry Judge Effect, derived from a study, reveals that judges tend to issue harsher sentences just before lunch, presumably influenced by hunger Danziger et al. (2011). This implies that physiological states, like hunger, may affect rational decision-making. Nevertheless, a subsequent study Weinshall-Margel and Shapard (2011) suggests that the observed peak in favorable decisions after a meal break is likely an artifact of case presentation order, considering anticipated outcomes and duration. It does not serve as evidence that physiological states, such as hunger, impact judges’ rational decision-making.
Appendix I National Judge
In ECtHR, a national judge refers to the judge appointed from the respondent country of the case which was brought before the court. Like all other judges, national judges maintain independence and do not act as representatives of their respective governments. However, as previously noted in studies by legal schorlars, national judges have been observed to dissent more frequently in cases finding a violation of the Convention compared to their non-national counterparts Kuijer (1997); Helfer and Voeten (2020).Voeten 2008 provides evidence and explanations including for why the judges are considered policy-seekers and the answers is that judicial activism is driven by the political logic of European integration.
Appendix J Key Case
KeyCase: The ECtHR annually chooses a set of significant cases, known as “key cases”. These often deal with complex and novel legal issues. Given the absence of established legal standards for interpreting them, they often generate controversy and disagreement among judges.
Appendix K High Reputation Country
Previous legal and political science work indicates ECtHR Judges tend to have more split vote in cases where the defending country is regarded as High Reputation Country. Strong Democracies enjoy a high reputation in front of the judges while weak and new democracies only benefit from a low reputation implying that the ECHR should issue more demanding judgments against low reputation States than against high reputation States with the view not to damage its own reputation and ensure compliance with its decision Dothan (2014). We extracted the country information from the HUDOC metadata, namely Respondent State(s). Following in Chalkidis et al. 2022b, we group the countries according to the disproportion of violations between Eastern and Central European countries, and the rest of European Countries (Western European, Nordic, Mediterranean states).
Appendix L Vulnerability in ECtHR
The European Court of Human Rights (ECtHR) adapts convention standards to meet individual needs and thus to ensure effective human rights protection. Recognizing vulnerability is crucial for understanding unique needs and implementing targeted support systems to empower individuals in need. However, the concept of vulnerability remains undefined at the ECtHR, which can lead to divergent opinions on what qualifies as vulnerability from a legal point of view. Cases involving vulnerable applicants often deal with complex social-political issues related to the protection and rights of individuals who may face various challenges, such as victimization, migration, discrimination, reproductive health, unpopular views etc. The inclusion of vulnerability as a proxy variable is an example of how a human rights legal concept that is difficult to define can potentially give rise to diverging opinions on the bench. A comprehensive study of such concepts in ECtHR jurisprudence lies beyond the scope of this work. We direct the reader to Heri 2021 for systematic legal study and Xu et al. 2023a for comprehensive NLP research of Vulnerability in ECtHR.
Appendix M Ablation Experiment
finetune dataset mic-F1 mac-F1 hm-F1
ours 73.55 ± 0.62 68.75 ± 1.13 68.03 ± 1.29
Santosh et al. 2022 73.41 ± 2.5 67.74± 3.2 63.93± 1.7
Table 7: LegalBERT’s COC performance on test set with different finetune dataset. Our dataset with correction of metadata as mentioned in § 4
We also train hierachical LegalBERTs with the textified allegation information on the original dataset from Santosh et al. (2022) without corrected metadata. Results in Tab 7 show that model achieve better performance when fine-tuned on our dataset with corrected metadata.
Appendix N Rationale for Avoiding Implementation of Uncertainty Evaluation Methods
Within the deep learning community, uncertainty is often classified into two types: aleatoric and epistemic uncertainty. Epistemic (or model) uncertainty arises from a lack of knowledge about the best model, often exacerbated by out-of-distribution examples. Aleatoric uncertainty, on the other hand, stems from inherent ambiguity and can be considered as the variability in experiment outcomes Houlsby et al. (2011); Gal et al. (2016). As SV is due to inherent disagreement among judges, we regard it as aleatoic uncertainty.
Aleatoric uncertainty is typically quantified through metrics such as Entropy Gal et al. (2016) and Softmax Response (SR) Geifman and El-Yaniv (2017), as recognized in prior works Malinin and Gales (2018). However, studies show that these methods are based on predictive entropy, and actually measure total uncertainty, combining both epistemic and aleatoric uncertainty Van Amersfoort et al. (2020). Only when we have prior knowledge that either aleatoric or epistemic uncertainty is low , can we use predictive entropy as a suitable measure for the other type Mukhoti et al. (2023).
Given our small, label-imbalanced, and temporally shifted dataset, we choose to refrain from utilizing these methods to directly measure the aleatoric uncertainty of our cases. Instead, we opt to assess difficulty using Predictive Variance of Information (PVI), which measures predictive entropy and can be considered an equivalent for total uncertainty.
Figure 7: Hierarchical Classification Variant
Appendix O Details of the Models
O.1 Architecture of the Hierarchical Model
For the hierarchical variant of pre-trained BERT models, We use a greedy input packing strategy where we merge multiple paragraphs into one packet until it reaches the maximum of 512 tokens. We independently encode each packet of the input text using the pretrained model and obtain representations ($h_{[CLS]}$) for each packet. Then we apply a non-pretrained transformer encoder to make the packet representations context-aware. Fig 7 illustrates the detailed architecture of the hierarchical model.
O.2 Implementation Details
We use BERT "bert-base-uncased" Kenton and Toutanova (2019), and LegalBERT "nlpaueb/legal-bert-base-uncased" Chalkidis et al. (2020) from the Transformers Hub Wolf et al. (2020) as our backbone model.
Hyperparameter & Overfitting Measures: For the Hierarchical models, we employ a maximum sentence length of 128 and document length (number of sentences) of 80. The dropout rate in all layers is 0.1. We follow the hyperparameters from Chalkidis et al. 2022a with a batch size of 8 and learning_rate of 3e-3. We train models with the Adam optimizer up to 10 epochs. We use PyTorch Paszke et al. (2019) 2.0.1.
Appendix P Soft Loss Function
The formulation of the soft loss function is represented as
$-\sum_{i=1}^{n}\sum_{c}p_{hum}\left(y_{i}\mid x_{i}\right)\log p_{\theta}\left(y_{i}=c\mid x_{i}\right)$, where we compute $p_{\text{hum }}$ by applying a standard normalization function to the judges’ votes for each pair following Peterson et al. 2019.
Figure 8: LegalBERT seed 1 on the SV cases in test set. Row 1: the distribution over instance-based absolute errors between probabilities for class 1 of the model vs human vote distributions. Row 2: the distribution over instance-based DistCE of mdodels. Row 3: the distribution over errors between probabilities for class 1 of the model vs human vote distributions.
Appendix Q Expected Calibration Error
Expected Calibration Error (ECE) is a measure of the difference between the predicted probabilities assigned by a model and the actual accuracy of those predictions. ECE is typically calculated by dividing the predicted probability space into a fixed number $M$ of intervals (or bins) and then computing the average absolute difference between the predicted and observed probabilities within each bin $B_{m}$ as follows:
$\mathrm{ECE}=\sum_{m=1}^{M}\frac{\left|B_{m}\right|}{N}\left|\operatorname{acc}\left(B_{m}\right)-\operatorname{conf}\left(B_{m}\right)\right|$ (1)
A lower ECE indicates better calibration, suggesting that the model’s predicted probabilities are more accurate reflections of the true likelihood. See Appendix for more about ECE.
Appendix R More Calibration Results
Tab 8 presents the calibration experiment results on both dev and test set. Soft-training improves the model’s COC performance (hm-F1) on SV cases for the development set, but lowers it for the test set. Two contributing factors account for this discrepancy: i) SV instances constitute less than 10% of the total dataset, potentially limiting the generalizability of improvements. ii) Given the temporal split of the dataset, distribution shifts in case-law may occur. For instance, SV cases could stabilize after a key legal decision. In other words, something that is legally controversial and uncertain in the training set time period may be settled law in the development and test dataset partitions. Our models may have overfit to the training set period. Fig 8 shows more figures comparing human uncertainty to model uncertainty.
[…]
Table 8: COC performance (hm-F1), confidence-calibration (ECE), and human-calibration (DistCE) performance with std($\pm$) in dev and test set. "/": COC finetuned models on § 6.1; "soft": models fine-tuned with soft-loss; "TS": model after Temperature Scaling. We choose the temperature $t$ by searching a range of possible values for $t$ on the dev set. We noted that the chosen $t$ across three random seeds were consistently overly aggressive, with values of 5.5, 5.8, and 5.5. Calibration results with std; Results shown over 3 random seeds.